A satellite revolves around the earth at a height h1 = 2Re. Calculate the height h2 of the satellite from the surface of the earth at which the speed of the satellite is 3/2v1 (in a radical), where v1 is the speed of the satellite at the height h1.
1
Expert's answer
2020-02-27T10:12:43-0500

From the conservation of energy:


"0.5mv_1^2-\frac{GmM}{2R_e}=0.5m(1.5v_1)^2-\frac{GmM}{h_2}"

"GM\left(\frac{1}{h_2}-\frac{1}{2R_e}\right)=0.625v_1^2"

"\frac{1}{h_2}=\frac{1}{2R_e}+\frac{0.625v_1^2}{GM}"

"h_2=\left(\frac{1}{2R_e}+\frac{0.625v_1^2}{GM}\right)^{-1}"